# Grammar Induction from Visual, Speech and Text

### v1

@@ 3 Methodology
encoder (dubbed VaTiora) framework based on the Diora model [8]. As shown in Fig. 3, VaTiora performs grammar induction with two key modules: the multimodal feature extraction and the inside-outside recursive autoencoding.
@@ 3.1 Feature Extraction
Textual Features. For text inputs, we follow the conventional GI methods, taking pre-trained word embeddings ELMo [40] to obtain the textual representations $\bm{W}=\{\bm{w}_{1},...,\bm{w}_{n}\}$ from the input $X$.
@@
For visual feature extraction, given the input image $I$, we follow [54] to adopt an external object detector (Faster-RCNN [10]) to extract a sequence of the object $\{o_{1},...,o_{M}\}$ along with their object features $\{\bm{v}_{1},...,\bm{v}_{M}\}$. Besides, each object $o_{m}$ has a bounding box $b_{m}$, which is encoded by an embedding layer to obtain the location features $\bm{l}_{m}$. We further consider the pair features of visual regions to determine the likelihood of object pairs forming a larger region. Concretely, we maintain a pair relevance matrix $\bm{M}\in\mathbb{R}^{C\times C}$ for all categories of object pairs in the dataset based on the co-occurrence probability, where $C$ is the total number of object categories. The calculation can be defined as:
$\bm{M}_{i,j}=\frac{\text{Count}(i,j)}{\sum_{l\in C}(\text{Count}(i,l)+\text{Count}(l,j))}\,,$ (2)
where $\text{Count}(i,j)$ means the count that objects with categories $i$ and $j$ detected in the same image within the entire dataset.
Speech Features. For speech features, we leverage a word discovery model VG-HuBert [38] to segment raw speech into clips $\{c_{1},...,c_{T}\}$, and obtain the clip representations $\{r_{1},...,r_{T}\}$. Note that each clip corresponds to a single word in the sentence when text is provided as input, i.e., $T=n$. To take into account the structural information of speech signals, we leverage pitch detection and voice activity detection (VAD) to extract intonation and rhythm features. First, we adopt the robust algorithm for pitch tracking (RAPT) [51], extracting the $f0$ value for each speech clip $c_{i}$, and encode it to a high dimension representation:
@@
where Avg denotes average operation, Round means rounding to an integer. $f0_{i}$ means the average $f0$ of frames in clip $c_{i}$. $\hat{\bm{p}}_{i}$ represents the pitch feature and will be used to enhance span representation. Moreover, we use the robust voice activity detection (rVAD) algorithm [52] to detect voice activity time $at_{i}$ and non-voice time $ut_{i}$ for each clip $c_{i}$.
@@ 3.2 Inside-Outside Recursive Autoencoder
Similar to Diora, VaTiora operates as an autoencoder, encoding the sentence into span representations and then reconstructing the bottom words of the constituent tree, i.e., the terminal nodes. The encoding process involves mapping the input words, denoted as $X$, to a latent constituent tree. To efficiently explore all valid trees, a dynamic programming-based inside-outside algorithm [33] is employed. Notably, VaTiora incorporates external multi-modal cues during the inside-outside pass, thereby extending the functionality of the Diora framework. Fig. 3 provides an overview of the interaction of features and the inside-outside pass.
@@
$\{\bm{h}_{i,i}^{in}\}=\text{CrossAttn}(\bm{W},\bm{R})\,.$ (4)
Similar to Diora, the VaTiora maintains a $N\times N$ chart $\bm{T}$ to store intermediate span vectors and scores, i.e., $\bm{h}_{i,j}^{in}$, $\bm{h}_{i,j}^{out}$ and $s_{i,j}^{in}$, $s_{i,j}^{out}$, for inside and outside representation respectively. We further fuse the visual and speech features into span vector $\bm{h}_{i,j}^{in}$ and score $s_{i,j}^{in}$. First, for each span $(i,j)$ and its decomposition $(i,k,j)$, we enhance $\bm{h}_{i,j}^{in}$ by:
$\begin{split}\hat{\bm{h}}_{i,j}^{in}&=\bm{h}_{i,j}^{in}+\gamma\bm{u}_{i,j}+\lambda\hat{\bm{p}}_{i,j},\\ \bm{h}_{i,j,k}^{in}&=\text{MLP}(\hat{\bm{h}}_{i,k}^{in};\hat{\bm{h}}_{k+1,j}^{in}),\end{split}$ (5)
where $1\leq i<k<j\leq N$, and (;)(;)( ; ) denotes the concatenation. $\gamma$ and $\lambda$ are the fusion weights. MLP is the composition function to merge two spans (see Diora [8]). Then $\bm{h}_{i,j}^{in}$ could be weighted sum of $\bm{h}_{i,j,k}^{in}$. $\bm{u}_{i,j}$ and $\bm{p}_{i,j}$ are the visual and pitch feature, obtained by:
$\begin{split}\bm{u}_{i,j}&=\sum_{m}^{M}(\bm{v}_{m}\oplus\bm{l}_{m})\cdot\text{Softmax}(\bm{v}_{m}^{\top}\bm{h}_{i,j}^{in}),\\ \hat{\bm{p}}_{i,j}&=\mathop{\text{AvgPool}}_{k\in[i,j]}(\hat{\bm{p}}_{k}),\end{split}$ (6)
where $\bm{v}_{m}$ is the object-level region features and $\bm{l}_{m}$ is the embedding of the bouding box coordinates $b_{m}$. $\bm{p}_{k}$ is the pitch feature obtained in Eq. (3). $\oplus$ denotes element-wise plus.
@@
$\begin{split}r_{i,j,k}=\sum_{o_{m},o_{n}}\text{Attn}_{i,k,o_{m}}&\cdot\text{Attn}_{k+1,j,o_{n}}\cdot\bm{M}_{\text{Cat}(o_{m}),\text{Cat}(o_{n})},\\ \text{Attn}_{i,k,o_{m}}&=\text{Softmax}(v_{m}^{\top}\bm{h}_{i,k}^{in}),\\ \text{Attn}_{k+1,j,o_{n}}&=\text{Softmax}(v_{n}^{\top}\bm{h}_{k+1,j}^{in})\end{split}$ (7)
@@
$a_{i,j,k}=-\text{Norm}(\text{Max}(ut_{t})/\text{Avg}(at_{t})),\;t\in[i,j)$ (8)
@@
$s_{i,j}^{in}=\sum_{k}s_{i,j,k}^{in}\text{Norm}(s_{i,j,k}^{in}).$ (10)
Inside-Outside Pass. The enhanced span vector $\bm{h}_{i,j}^{in}$ and score $s_{i,j}^{in}$ are then utilized to perform an inside-outside pass similar to Diora. In the inside pass, each span $(i,j)$ computes inside vector $\bm{h}_{i,j}^{in}$ and score $s_{i,j}^{in}$ by weighted summing all possible substructures. Similarly, the outside pass reverses this process by computing $\bm{h}_{i,j}^{out}$ and $s_{i,j}^{out}$ from top to bottom, where the bottom-most vector $\bm{h}_{i,i}^{out}$ is used to reconstruct the words. Finally, the span score is calculated as $q(i,j)=s_{i,j}^{in}\cdot_{i,j}^{out}/s_{1,n}^{in}$ to measure how likely the span $(i,j)$ exists. More technical details on the inside and outside passing in VaTiora are provided in Appendix §C.
@@ 3.3 Overall Training
Structure Reconstruction and Contrastive Learning. Following [54, 8], we adopt structure reconstruction and contrastive learning For reconstruction, a self-supervised blank-filling objective is defined as:
@@
$\begin{split}l_{span}(I,i,j)=&\max\{0,d(I,(i,j)^{\prime})-d(I,(i,j))+\epsilon\}\\ &+\max\{0,d(I^{\prime},(i,j))-d(I,(i,j))+\epsilon\}\,,\\ d(I,(i,j))=&sim((i,j),I)\times q(i,j),i\neq j,\\ sim(i,j,I)=&\max_{m\in[0,M]}\{\bm{v}_{m}^{\top}(\bm{h}_{i,j}^{in}+\bm{h}_{i,j}^{out})\},\end{split}$ (12)
@@
where $\epsilon$ is the positive margin. For bottom-most $(i,i)$, e.g., the word $w_{i}$ (or the speech clip $c_{i}$ in textless setting), we compute:
@@
$\mathcal{L}_{rep}=\text{Cos}(\bm{r}_{i},\bm{w}_{i})+\text{Cos}(\text{Avg}(\{\bm{r}_{i}\}),\text{Avg}(\{\bm{v}_{m}\})),$ (15)
where Cos means cosine similarity. $\text{Avg}(\{\bm{r}_{i}\})$, $\text{Avg}(\{\bm{v}_{m}\})$ means average of the clip and object vectors, representing the whole speech and image feature. In textless setting, there are only $\text{Cos}(\text{Avg}(\{\bm{r}_{i}\})$ and $\text{Avg}(\{\bm{v}_{m}\}))$. The whole training loss will be:
@@
In the inference stage, we follow the conventional GI method to predict the tree with the maximum inside scores via the CKY algorithm [22]. The details are shown in Appendix §D.
4 Novel Dataset
VAT-GI can be evaluated on the SpokenCOCO dataset [17], which contains approximately 600,000 speech recordings of MSCOCO image captions [^2] https://cocodataset.org/. Nevertheless, SpokenCOCO can fall prey to limitations in terms of its visual scenario styles, short captions, and monotonous speech tone, resulting in inferior generalization performance of VAT-GI parser across diverse language environments. Thus, we introduce a more challenging test set, namely SpokenStory. The dataset is built by extending the image caption data with speech records. We comprises 1,000 images extracted from OpenImages [31] across various scenarios, and accompanied by corresponding caption annotations from [43]. OpenImages is widely used for computer vision research and applications, which contains over 9 million labeled images sourced from the web and covers more than 6,000 topics and categories. Each image is annotated with labels that describe the objects or concepts present in the image. We adopt the annotations of Localized Narratives [43] for aligned captions, and we manually record speech for each caption by human speakers, who are recruited in our research team, producing the aligned transcripts and speech audios. The average sentence length of SpokenStory is 20 words, with an average speech duration of 5.01 seconds, roughly twice that of SpokenCOCO. The speech recordings encompass five distinct tone styles, i.e., “Monotonous”, “Excited”, “Tense”, “Peaceful” and “Eccentric”, with each characterized by specific intonation and rhythm. We do not suggest speakers how to express each emotion while letting them express all by their own comprehension. Overall, the new test set poses a more challenging setting, due to its longer sentence, complex constituents and diverse intonations and tones. Table 2 compares two datasets.
Table 2: Statistics of SpokenCOCO and SpokenStory. We follow [21] to split SpokenCOCO.

### v2

@@ 4 Framework of VaTiora
encoder (dubbed VaTiora) framework based on the Diora model Drozdov et al. (2019). As shown in Figure 3, VaTiora performs grammar induction with two key modules: the multimodal feature extraction and the inside-outside recursive autoencoding.
@@ 4.1 Feature Extraction
Textual Features. For text inputs, we follow the conventional GI methods, taking pre-trained word embeddings ELMo Peters et al. (2018) to obtain the textual representations $\bm{W}=\{\bm{w}_{1},...,\bm{w}_{n}\}$ from the input $X$. In practice, we can also adopt other pre-trained word embeddings, such as Glove, or use randomly initialized embedding.
@@
For visual feature extraction, given the input image $I$, we follow Wan et al. (2022) to adopt an external object detector (Faster-RCNN Girshick (2015)) to extract a sequence of the object $\{o_{1},...,o_{M}\}$ along with their object RoI features $\{\bm{v}_{1},...,\bm{v}_{M}\}$. Besides, each object $o_{m}$ has a bounding box $b_{m}$, which is encoded by an embedding layer to obtain the location features $\bm{l}_{m}$. Denote $b_{m}$ as:
$b_{m}=(x_{min},y_{min},x_{max},y_{max}).$ (2)
With the top-left as the origin point, we normalize it to:
$b^{\prime}_{m}=(x_{min}/w,y_{min}/h,x_{max}/w,y_{max}/h),$ (3)
$\bm{l}_{m}=\text{FC}(b_{m}).$ (4)
We further consider the pair features of visual regions to determine the likelihood of object pairs forming a larger region. Concretely, we maintain a pair relevance matrix $\bm{M}\in\mathbb{R}^{C\times C}$ for all categories of object pairs in the dataset based on the co-occurrence probability, where $C$ is the total number of object categories. Supposing there are $C$ categories of object label, and for every two object that belong to $c_{\alpha},c_{\beta}\in C$, their score is:
$\bm{M}_{\alpha,\beta}=\frac{\text{Co-Count}(\alpha,\beta)}{\sum_{\gamma\in C}(\text{Co-Count}(\alpha,\gamma)+\text{Co-Count}(\gamma,\beta))}\,,$ (5)
[…]
Speech Features. For speech features, we leverage a word discovery model VG-HuBert Peng and Harwath (2022) to segment raw speech into clips $\{c_{1},...,c_{T}\}$, and obtain the clip representations $\{r_{1},...,r_{T}\}$. In the full setting, we take the textual transcripts to force align the speech, while in the textless setting, we use VG-HuBert to perform unsupervised word discovery. Note that each clip corresponds to a single word in the sentence when text is provided as input, i.e., $T=n$. To take into account the structural information of speech signals, we leverage pitch detection and voice activity detection (VAD) to extract intonation and rhythm features. First, we adopt the robust algorithm for pitch tracking (RAPT) Talkin and Kleijn (1995), extracting the $f0$ value for each speech clip $c_{i}$, and encode it to a high dimension representation. Moreover, considering the pitch frequency constraint of normal human speech, we limit the extracted $f0$ value to the range of 50Hz to 500Hz Trollinger (2003). Then, we embed the rounded $f0$ (450 values) to the high-dimension representation, and use an LSTM model to capture the temporal features of pitch changing:
@@
where Avg denotes average operation, Round means rounding to an integer. $f0_{i}$ means the average $f0$ of frames in clip $c_{i}$. $\hat{\bm{p}}_{i}$ represents the pitch feature and will be used to enhance span representation. At last, we represent the rhythm feature of speech. The voice activity time can be extracted by the robust voice activity detection (rVAD) method Tan et al. (2020). For each clip, VAD outputs the number of speech frames and the non-speech frames. We sum the frame length of speech frames and non-speech frames as voice activity time $at_{i}$ and non-voice time $wt_{i}$ for each clip $c_{i}$.
@@ 4.2 Inside-Outside Recursive Autoencoder
Similar to Diora, VaTiora operates as an autoencoder, encoding the sentence into span representations and then reconstructing the bottom words of the constituent tree, i.e., the terminal nodes. The encoding process involves mapping the input words, denoted as $X$, to a latent constituent tree. To efficiently explore all valid trees, a dynamic programming-based inside-outside algorithm Lari and Young (1990) is employed. Notably, VaTiora incorporates external multi-modal cues during the inside-outside pass, thereby extending the functionality of the Diora framework. Figure 3 provides an overview of the interaction of features and the inside-outside pass.
@@
$\begin{split}&\{\bm{h}_{i,i}^{in}\}=\text{CrossAttn}(\bm{W},\bm{R})\,,\\ &\text{CrossAttn}(\bm{W},\bm{R})=\text{Softmax}((\bm{QW})(\bm{KR}))\bm{VR},\end{split}$ (7)
where the $\bm{Q}$, $\bm{K}$, $\bm{V}$ are training parameters. The bottom-most node span score $s_{i,i}^{in}$ is initialized with 0. Similar to Diora, the VaTiora maintains a $N\times N$ chart $\bm{T}$ to store intermediate span vectors and scores, i.e., $\bm{h}_{i,j}^{in}$, $\bm{h}_{i,j}^{out}$ and $s_{i,j}^{in}$, $s_{i,j}^{out}$, for inside and outside representation respectively. We further fuse the visual and speech features into span vector $\bm{h}_{i,j}^{in}$ and score $s_{i,j}^{in}$. First, for each span $(i,j)$ and its decomposition $(i,k,j)$, we enhance $\bm{h}_{i,j}^{in}$ by:
$\begin{split}&\hat{\bm{h}}_{i,k}^{in}=\bm{h}_{i,k}^{in}+\gamma\bm{u}_{i,k}+\lambda\hat{\bm{p}}_{i,j},\\ &\bm{h}_{i,j,k}^{in}=\text{MLP}(\hat{\bm{h}}_{i,k}^{in};\hat{\bm{h}}_{k+1,j}^{in}),\\ &\bm{h}_{i,j}^{in}=\sum_{k}\bm{h}_{i,j,k}^{in}\cdot\mathop{\text{Softmax}}_{k}(s_{i,j,k}^{in}),\end{split}$ (8)
where $1\leq i<k<j\leq N$, and (;)(;)( ; ) denotes the concatenation. $\gamma$ and $\lambda$ are the fusion weights. The MLP (Multi-Layer Perceptron) is the network contains multiple linear layers, i.e., $\bm{h}_{l+1}=\sigma(\bm{W}_{l}\bm{h}_{l}+\bm{b}_{l})$, where $\sigma$ is an activation function, $\bm{W}_{l}$ and $\bm{b}_{l}$ are training parameters. We adopt a two-layer MLP here, which is used as the composition function to merge two spans (see Diora Drozdov et al. (2019)). The $s_{i,j,k}^{in}$ is the decomposition score calculated during the inside pass (Eq. 12). Then $\bm{h}_{i,j}^{in}$ could be weighted sum of $\bm{h}_{i,j,k}^{in}$. $\bm{u}_{i,j}$ and $\bm{p}_{i,j}$ are the visual and pitch feature, obtained by:
$\begin{split}\bm{u}_{i,j}&=\sum_{m}^{M}(\bm{v}_{m}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}\bm{l}_{m})\cdot\text{Softmax}(\bm{v}_{m}^{\top}\bm{h}_{i,j}^{in}),\\ \hat{\bm{p}}_{i,j}&=\mathop{\text{AvgPool}}_{k\in[i,j]}(\hat{\bm{p}}_{k}),\end{split}$ (9)
where $\bm{v}_{m}$ is the object-level region features and $\bm{l}_{m}$ is the embedding of the bouding box coordinates $b_{m}$. $\hat{\bm{p}}_{k}$ is the pitch feature obtained in Eq. (6). AvgPool($\cdot$) is the vector averaging operation.
Figure 4: Illustration of pair feature and voice activity feature.
@@
$\begin{split}r_{i,j,k}=\sum_{o_{m},o_{n}}\text{Attn}_{i,k,o_{m}}&\cdot\text{Attn}_{k+1,j,o_{n}}\cdot\bm{M}_{\text{Cat}(o_{m}),\text{Cat}(o_{n})},\\ \text{Attn}_{i,k,o_{m}}&=\text{Softmax}(\bm{v}_{m}^{\top}\bm{h}_{i,k}^{in}),\\ \text{Attn}_{k+1,j,o_{n}}&=\text{Softmax}(\bm{v}_{n}^{\top}\bm{h}_{k+1,j}^{in})\end{split}$ (10)
@@
$a_{i,j,k}=-\text{Sigmoid}(\text{Max}(ut_{t})/\text{Avg}(at_{t})),\;t\in[i,j)$ (11)
@@
$s_{i,j}^{in}=\sum_{k}s_{i,j,k}^{in}\mathop{\text{Softmax}}_{k}(s_{i,j,k}^{in}).$ (13)
Figure 4 visually illustrates the contributions of these features.
Figure 5: The process of inside and outside pass, respectively.
Inside-Outside Pass. The enhanced span vector $\bm{h}_{i,j}^{in}$ and score $s_{i,j}^{in}$ are then utilized to perform an inside-outside pass similar to Diora. The process is based on the inside-outside algorithm Baker (1979), which is used for the derivation process of probabilistic context-free grammar. This method produces representations for all internal nodes via recursively filling a chart, where each cell represents a node in the latent tree.
The inside pass of this method recursively compresses the input sequence, at each step inputting the vector representations of the two children into a composition function (i.e. Equation 8) that outputs an inside vector representation of the parent. This process continues up to the root of the tree, eventually yielding a single vector representing the entire sentence (Figure 5 left part). In the inside pass, each span $(i,j)$ computes inside vector $\bm{h}_{i,j}^{in}$ and score $s_{i,j}^{in}$ by weighted summing all possible substructures.
Following that, the outside pass computes the outside presentations via a top-down process (Figure 5 right part). The outside representations are encoded by looking at only the context of a given sub-tree. The root node of the outside chart is learned as a bias. Descendant cells are predicted using a disambiguation over the possible outside contexts. Each component of the context consists of a sibling cell from the inside chart and a parent cell from the outside chart. In the outside pass, we compute $\bm{h}_{i,j}^{out}$ and $s_{i,j}^{out}$ from top to bottom, where the bottom-most vector $\bm{h}_{i,i}^{out}$ is used to reconstruct the words. For a span $(i,j)$ and a $k$ out of $(i,j)$:
Similarly to the inside pass, the outside span representation $\bm{h}_{i,j}^{out}$ and score $s_{i,j}^{out}$ is:
$\begin{split}\bm{h}_{i,j}^{out}&=\sum_{k}\bm{h}_{i,j,k}^{out}\cdot\mathop{\text{Softmax}}_{k}(s_{i,j,k}^{out}),\\ s_{i,j}^{out}&=\sum_{k}s_{i,j,k}^{out}\cdot\mathop{\text{Softmax}}_{k}(s_{i,j,k}^{out}),\end{split}$ (16)
Finally, the span score is calculated as $q(i,j)=s_{i,j}^{in}\cdot s_{i,j}^{out}/s_{1,n}^{in}$ to measure how likely the span $(i,j)$ exists.
@@ 4.3 Overall Training
Structure Reconstruction and Contrastive Learning. Following Wan et al. (2022); Drozdov et al. (2019), we adopt structure reconstruction and contrastive learning For reconstruction, a self-supervised blank-filling objective is defined as:
@@
$\begin{split}&l_{span}(I,i,j)=\max\{0,d(I,(i,j)^{\prime})-d(I,(i,j))+\epsilon\}\\ &\quad\quad\quad\quad+\max\{0,d(I^{\prime},(i,j))-d(I,(i,j))+\epsilon\}\,,\\ &d(I,(i,j))=sim((i,j),I)\times q(i,j),i\neq j,\\ &sim(i,j,I)=\max_{m\in[0,M]}\{\bm{v}_{m}^{\top}(\bm{h}_{i,j}^{in}+\bm{h}_{i,j}^{out})\},\end{split}$ (18)
@@
where $\epsilon$ is the positive margin. $q(i,j)$ is the span score as mentioned above ($q(i,j)=s_{i,j}^{in}\cdot_{i,j}^{out}/s_{1,n}^{in}$). For bottom-most $(i,i)$, e.g., the word $w_{i}$ (or the speech clip $c_{i}$ in textless setting), we compute:
@@
where Cos means cosine similarity. $\text{Avg}(\{\bm{r}_{i}\})$, $\text{Avg}(\{\bm{v}_{m}\})$ means average of the clip and object vectors, representing the whole speech and image feature. In textless setting, there are only $\text{Cos}(\text{Avg}(\{\bm{r}_{i}\}),\text{Avg}(\{\bm{v}_{m}\}))$. The whole training loss will be:
@@
In the inference stage, we follow the conventional GI method to predict the tree with the maximum inside scores via the CKY algorithm.
Figure 7: Comparison of captions between our SpokenStory and SpokenCOCO.
Figure 8: Visualized waveform for the sentence “A woman scanning a tour map with her smartphone” in the five emotions.
Table 2: Statistics of SpokenCOCO and SpokenStory. We follow Karpathy and Fei-Fei (2017) to split SpokenCOCO. “Avg. Sppech Len.” means average speech length. “Agb. Sent. Len.” means average sentence length.
@@
5 Novel Dataset for VAT-GI
VAT-GI can be evaluated on the SpokenCOCO dataset Hsu et al. (2021), which contains approximately 600,000 speech recordings of MSCOCO image captions [^2] https://cocodataset.org/. Nevertheless, SpokenCOCO can fall prey to limitations in terms of its visual scenario styles, short captions, and monotonous speech tone, resulting in inferior generalization performance of VAT-GI parser across diverse language environments. Thus, we introduce a more challenging test set, namely SpokenStory. The dataset is built by extending the image caption data with speech records. We comprises 1,000 images extracted from OpenImages Kuznetsova et al. (2020) across various scenarios, and accompanied by corresponding caption annotations from Pont-Tuset et al. (2020). OpenImages is widely used for computer vision research and applications, which contains over 9 million labeled images sourced from the web and covers more than 6,000 topics and categories. Each image is annotated with labels that describe the objects or concepts present in the image.
We adopt the annotations of Localized Narratives Pont-Tuset et al. (2020) for aligned captions, and we manually record speech for each caption by human speakers, who are recruited in our research team, producing the aligned transcripts and speech audios. The average sentence length of SpokenStory is 20 words, with an average speech duration of 5.01 seconds, roughly twice that of SpokenCOCO. Figure 7 shows the comparison of captions between our SpokenStory and SpokenCOCO, where the sentence in SpkenStory has longer constituents. The speech recordings encompass five distinct tone styles, i.e., “Monotonous”, “Excited”, “Tense”, “Peaceful” and “Eccentric”, with each characterized by specific intonation and rhythm. We do not suggest speakers how to express each emotion while letting them express all by their own comprehension. Figure 8 illustrates the waveform of five types of speech in the same sentence. We develop a tool for recording and transcript annotation. The speakers first record the captions and the system will return a pseudo transcript via an off-the-shelf speech segmentation model. After recording, the speakers are asked to adjust the transcript to check the word boundaries. Then the recording and the transcripts will be saved into the database. Overall, the new test set poses a more challenging setting, due to its longer sentence, complex constituents and diverse intonations and tones. Table 2 compares two datasets.